### v2

# Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations

Abstract
The advancement of AI systems for mental health support is hindered by limited access to therapeutic conversation data, particularly for trauma treatment. We present Thousand Voices of Trauma, a synthetic benchmark dataset of 3,000 therapy conversations based on Prolonged Exposure therapy protocols for Post-traumatic Stress Disorder (PTSD). The dataset comprises 500 unique cases, each explored through six conversational perspectives that mirror the progression of therapy from initial anxiety to peak distress to emotional processing. We incorporated diverse demographic profiles (ages 18-80, M=49.3, 49.4% male, 44.4% female, 6.2% non-binary), 20 trauma types, and 10 trauma-related behaviors using deterministic and probabilistic generation methods. Analysis reveals realistic distributions of trauma types (witnessing violence 10.6%, bullying 10.2%) and symptoms (nightmares 23.4%, substance abuse 20.8%). Clinical experts validated the dataset’s therapeutic fidelity, highlighting its emotional depth while suggesting refinements for greater authenticity. We also developed an emotional trajectory benchmark with standardized metrics for evaluating model responses. This privacy-preserving dataset addresses critical gaps in trauma-focused mental health data, offering a valuable resource for advancing both patient-facing applications and clinician training tools.
Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations
Suhas BN1  Dominik Mattioli1  Saeed Abdullah1 Rosa I. Arriaga2  Chris W. Wiese3  Andrew M. Sherrill4 1College of Information Sciences and Technology, Penn State University, USA 2School of Interactive Computing, Georgia Tech, USA 3School of Psychology, Georgia Tech, USA 4Department of Psychiatry and Behavioral Sciences, Emory University, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
1 Introduction
The intersection of mental health care and artificial intelligence presents unprecedented opportunities alongside significant challenges. AI system development faces particular obstacles in trauma-focused therapy due to the sensitive nature of patient experiences and strict privacy regulations, which make the collection of real-world data extremely challenging (Mathur et al., 2022).
Moreover, existing datasets frequently lack the diversity and clinical depth needed to train robust AI systems capable of serving diverse populations effectively Marr (2023). Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD) (Foa et al., 2007)—offers a structured therapeutic approach that could especially benefit from AI support.
However, current mental health conversation datasets are often too small (Zantvoort et al., 2024), lack demographic diversity (Sadeh-Sharvit, 2024), and do not capture the nuanced progression of trauma-focused therapy sessions (Mathur et al., 2022).
To address these gaps, we introduce Thousand Voices of Trauma, a synthetic benchmark dataset comprising 500 clinical sessions. Each session is divided into six distinct therapeutic phases aligned with core evaluation criteria of PE therapy. Each phase contains multiple therapist-client exchanges, which can be analyzed either independently or as part of the complete session flow, resulting in 3000 distinct clinical conversations. These phases, as defined by Foa et al. (2007), correspond to specific stages within the PE therapy session: a) Orientation to Imaginal Exposure, b) Imaginal Exposure Duration, c) Monitoring SUDS Ratings, d) Reinforcing Comments, e) Eliciting Thoughts and Feelings, and f) Processing the Imaginal.
These phases mirror the typical session progression—from initial anxiety, through peak distress during imaginal exposure, to the gradual reduction of distress through reinforcement and processing. Its diversity encompasses a wide range of demographic profiles, trauma types, and associated behaviors, designed to reflect varied real-world clinical presentations. This structured, diverse dataset offers scalable opportunities for AI systems to assist mental health professionals in trauma-focused therapy. By emphasizing diverse populations, Thousand Voices of Trauma represents a meaningful step toward more effective, personalized, and ethically guided mental health care.
This dataset also addresses other real-world limitations that often hinder mental health research and safe AI model development. For example, while privacy concerns often restrict data access, synthetic data can circumvent typical ethical and legal barriers. It also overcomes other common issues with real-world data, such as incompleteness, inconsistency, and small sample sizes—especially among minority groups.
By balancing representation across diverse populations, trauma types, and racial or ethnic minorities, the dataset helps mitigate inherent biases. For instance, the NIMH reported in 2021 that 14.5 million U.S. adults (5.7%) experienced severe major depressive episodes, with higher rates among females (10.3%) than males (6.2%) and the highest prevalence among those aged 18–25 (18.6%). Synthetic data can compensate for such imbalances, enhancing model training and analysis.
1.1 Key Contributions
  1. 1.
    Scale and Diversity: Our dataset is one of the largest collections of synthetic trauma-focused therapy conversations, encompassing varied demographic profiles across age, gender, ethnicity, and cultural backgrounds. The use of synthetic data addresses ethical concerns related to patient privacy and consent, while its diversity supports the development of more inclusive and culturally competent AI systems.
  2. 2.
    Clinical Depth: Each conversation is grounded in evidence-based PE therapy principles and incorporates 20 distinct trauma types, 10 trauma-related behaviors, and 5 co-occurring conditions. This clinical specificity renders the dataset particularly valuable for specialized therapeutic applications, including clinician training.
  3. 3.
    Structured Evaluation Framework: The six-conversation structure per session enables a comprehensive examination of different interaction trajectories, spanning from the initial assessment to trauma processing and progress evaluation.
  4. 4.
    Baseline: We establish a baseline dataset for training AI systems in trauma-focused therapy, facilitating standardized comparisons across different approaches.
The remainder of this paper is structured as follows: Section 2 reviews related work in mental health. Section 3 describes the datasets and details the synthetic data generation methodology. Section 4 presents the expert evaluation. Section 5 introduces the benchmark setup and evaluation method. Sections 69 discuss future directions, data availability, ethical considerations, and limitations.
2 Related Work
Prolonged Exposure (PE) therapy, an evidence-based treatment for PTSD, relies on structured exposure to trauma-related narratives (Watkins et al., 2018). However, there is a lack of trained professionals who can provide PE therapy (Rauch et al., 2017, 2023). As a result, there is an urgent need for AI applications to support PE therapy delivery and training. This underscores the need for clinically valid and diverse datasets for AI development and evaluation. Large-scale language models (LLMs), like the GPT series, have shown potential in generating synthetic datasets that mimic human-like text, addressing challenges such as data scarcity and privacy concerns (Hämäläinen et al., 2023; Giuffrè and Shung, 2023; Zhu, 2024; Li et al., 2024). However, for applications like PE therapy—which require alignment with trauma-focused frameworks, diverse demographic representation, and strict ethical safeguards—current research still lacks tailored solutions.
Synthetic datasets show promise in mental health applications, with studies exploring LLM-based data generation to address data scarcity. Wu et al. (Wu et al., 2023, 2024) introduced zero-shot and few-shot learning frameworks to augment PTSD diagnostic datasets, producing synthetic transcripts that outperform baselines. The latter work used role-prompting and structured prompts to create realistic synthetic clinical interviews.
However, these approaches have not been specifically adapted to the structured narratives essential for PE therapy. Efforts to enhance demographic diversity in synthetic datasets are growing. Mori et al. (2024) and Lozoya et al. (2023) examined how synthetic data reflects demographic variation, highlighting biases in LLM outputs, especially regarding race and gender, and stressing fairness in mental health datasets. Techniques like patient vignette simulation (Denecke, 2024) and adaptive prompts for non-English contexts (e.g., SAPE for Spanish (Lozoya et al., 2024)) show early progress toward inclusivity. However, trauma-type diversity and PE-specific scenarios remain unexplored. Additionally, Chen et al. (2024) underscored the need for systematic benchmarking using metrics like F1-score, AUC, and balanced accuracy, but these tools have not yet been applied to datasets focused on diverse trauma types or PE therapy.
Privacy and ethics are central to synthetic dataset generation, with studies focusing on privacy-preserving methods. Recent works (Chuang et al., 2024a; Meoni et al., 2024; Suhas et al., 2023; BN and Abdullah, 2022; Chuang et al., 2024b) highlight privacy-preserving machine learning, protected health information (PHI) exclusion, and semantic filtering to maintain privacy compliance while preserving data utility. While these methods offer strong safeguards for general clinical use, ethical risks—such as generating harmful trauma narratives or victim-blaming—remain underexplored, especially in sensitive contexts like trauma-focused therapies such as PE.
While synthetic dataset generation using LLMs (Wu et al., 2023, 2024; Xu et al., 2024) has advanced in addressing general clinical challenges, gaps in the literature remain for PE therapy. These include the lack of trauma-type diversity, limited demographic inclusivity evaluation, insufficient alignment with frameworks like DSM-5 PTSD criteria, and underdeveloped ethical safeguards specific to trauma-focused contexts. Synthetic datasets can also help mitigate representation biases in AI models for mental healthcare delivery.
For example, American Psychiatric Association (2013) reported that non-Hispanic White adults (25.0%) were more likely to receive mental health services than non-Hispanic Black (18.3%), Hispanic (17.3%), and Asian (13.9%) adults. Including such underrepresented groups in synthetic datasets might partially address the training data gap. This paper seeks to bridge these gaps by exploring the existing knowledge base and identifying pathways to tailor synthetic data generation for PE therapy applications.
Figure 1: The figure presents the demographic distribution of generated participant profiles across four categories: gender, age, ethnicity, and relationship status. The gender distribution shows a nearly balanced representation of males (247) and females (222), with a smaller group identifying as non-binary (31) mimicking real world gender distributions U.S. Census Bureau (2025). The Age Distribution spans from under 10 to over 90 years, with most participants between 30 and 70 years. The Ethnicity Distribution highlights diverse backgrounds, with the largest groups being Latin American (70), North American (67), South East Asian (64), South Asian (63), and European (63), followed by Oceania (62), Middle Eastern (59), and African (52) participants. The Relationship Status chart reveals that most participants are married (180) or single (149), while smaller numbers are in relationships (70), divorced (51), or widowed (50).
Figure 2: The figure presents the distribution of exhibited behaviors and types of trauma among participants. The Exhibited Behaviors chart shows that nightmares (117 participants) are the most common symptom, followed by substance abuse (104), compulsive behaviors (101), aggression (100), and hypervigilance (99). Other frequently reported behaviors include self-blame (97), withdrawal (95), flashbacks (95), avoidance (91), and dissociation (83). The Types of Trauma chart highlights that witnessing violence (53 participants) and bullying (51) are the most frequently reported traumatic experiences. Other significant trauma types include neglect (45), medical trauma (43), childhood trauma (42), and the loss of a loved one (39). Natural disasters, sexual abuse, physical abuse, and emotional abuse were each reported by 39 to 38 participants, while combat or war experiences and accidents were each cited by 36 participants.
3 Dataset
3.1 Dataset: Simulated Therapy Session Profiles
We generated a synthetic dataset of therapy session profiles to analyze trauma narratives and therapeutic interactions. See Appendix A for full prompt templates used to generate each therapeutic phase. Profiles include client demographics, therapist characteristics, and session variables such as trauma type and topic. A combination of deterministic and probabilistic methods ensured diversity and realistic representation of therapeutic scenarios.
3.2 Client Profile Generation
Client profiles included age, gender, relationship status, occupation, living situation, and ethnicity. Ages ranged from 18 to 80, divided into six groups: 18-30, 31-40, 41-50, 51-60, 61-70, 71-80. We assigned gender using weighted probabilities: 50% male, 49% female, and 1% non-binary U.S. Census Bureau (2025). Relationship status, occupation, and living situation were age-specific—for example, clients aged 20-30 were more likely to be “Single,” “Student,” and “With parents,” while those 60-70 were often “Widowed,” “Retired,” and “Alone.” A validation function ensured logical consistency. We randomly assigned ethnicity from eight global regions: South Asian, Middle Eastern, African, North American, Oceanian, European, South East Asian, and Latin American. We assigned co-occurring conditions with weighted probabilities Bilevicius et al. (2020); Jennifer et al. (2024); Hagiwara et al. (2022): None (25%), Anxiety (25%), Depression (30%), Substance Use Disorder (10%), and Chronic Pain (10%). We also assigned clients 1–3 trauma-related behaviors from ten options, including avoidance, hypervigilance, flashbacks, nightmares, self-blame, substance abuse, aggression, withdrawal, dissociation, and compulsive behaviors. The options represent a range of cognitive, emotional, and behavioral responses typically associated with trauma, which aligns with trauma-informed care principles (SAMHSA, 2015, 2014).
3.3 Therapist Profile Generation
We generated therapist profiles with ages ranging from 25 to 65, divided into four age groups: 25-34, 35-44, 45-54, and 55-65. We assigned therapist gender using the same weighted probabilities as client gender.
3.4 Therapy Context Generation
To create varied therapeutic contexts, we defined both a broad trauma type representing the general category of traumatic experience, and a more specific session topic providing the focus for the simulated interaction. We generated the therapy context by randomly selecting a trauma type from twelve possibilities (SAMHSA, 2015, 2014): physical abuse, emotional abuse, sexual abuse, neglect, natural disasters, accidents, combat or war experiences, loss of a loved one, witnessing violence, bullying, childhood trauma, and medical trauma. A session topic was also randomly selected from twenty possibilities, such as car accident trauma, workplace trauma, domestic violence, natural disaster experience, military combat experience, loss of a loved one, severe illness, divorce or relationship breakup, witnessing violence, racial trauma, and refugee experiences. This independent random selection of trauma type and session topic allows for a wide range of generated scenarios. While this methodology can result in pairings without an immediate thematic link (e.g., ‘natural disaster’ type with ‘workplace trauma’ topic), such combinations can also reflect the complexity observed in real-world therapy, where session discussions may focus on various life stressors influenced by, or co-existing with, the client’s primary trauma history.
3.5 Session Profile Assembly
Each complete session profile combined a validated client profile, a therapist profile, and the generated therapy context, including the trauma type and session topic.
3.6 Dataset Statistics
The dataset comprises 500 simulated participants (ages 18-80 years, M = 49.3). The gender distribution includes 247 male (49.4%), 222 female (44.4%), and 31 non-binary (6.2%) participants.
The ethnicity distribution is as follows: Latin American (70, 14.0%), North American (67, 13.4%), South East Asian (64, 12.8%), South Asian (63, 12.6%), European (63, 12.6%), Oceania (62, 12.4%), Middle Eastern (59, 11.8%), and African (52, 10.4%). Regarding relationship status, participants were predominantly married (180, 36.0%) or single (149, 29.8%), with others reporting being in a relationship (70, 14.0%), divorced (51, 10.2%), or widowed (50, 10.0%). See Figure 1 & 2 for more details.
Figure 3: The figures illustrate structure and language diversity in synthetic therapist-client dialogues. The Utterance Length Distribution (top) shows clients often speak at length (>50 words), while therapists’ responses are concise, reflecting the client-centered nature of therapy. The Vocabulary Diversity (bottom) reveals clients use ~24,000 unique words, far more than therapists (~5,000), likely due to personal narratives, whereas therapists maintain structured, reflective language.
Figure 4: The figure depicts conversation flow in synthetic dialogues, showing exchange lengths over time across three phases: Setup, Exposure, and Processing. In Setup, lengths remain stable (~40 to 45 words). Exposure sees a steady increase, peaking at ~60 words, indicating deeper engagement. Processing shows fluctuations, reflecting varying reflection and emotional processing. The shaded region represents variability across conversations.
Generated interactions exhibited various trauma-related behaviors (See Figure 2, 3 and 4 respectively), with nightmares being most prevalent (117, 23.4%), followed by substance abuse (104, 20.8%), compulsive behaviors (101, 20.2%), and aggression (100, 20.0%). Other common manifestations included hypervigilance (99, 19.8%), self-blame (97, 19.4%), withdrawal and flashbacks (95 each, 19.0%), avoidance (91, 18.2%), and dissociation (83, 16.6%). The types of trauma are diverse, with witnessing violence being most common (53, 10.6%), followed by bullying (51, 10.2%), neglect (45, 9.0%), medical trauma (43, 8.6%), and childhood trauma (42, 8.4%). Other reported traumas included loss of a loved one, natural disasters, sexual abuse, and physical abuse (39 each, 7.8%), emotional abuse (38, 7.6%), and combat or war experiences and accidents (36 each, 7.2%).
Figure 5: Stacked Bar charts illustrating therapist ratings (N=7) across four key dimensions of the synthetic PE therapy sessions—Content Depth, Perceived Value, Session Appropriateness, and Patient Engagement.
Figure 6: Violin plots illustrating the distribution of emotional intensity scores across six therapy stages scored by Claude Sonnet 3.5. Emotional intensity generally starts at moderate levels during orientation, peaks during exposure and monitoring, and then tapers off through reinforcing, reflection, and processing phases.
4 Expert Evaluation of Synthetic PE Therapy Sessions
To ensure clinical relevance and gain insights into the utility and limitations of our synthetic dataset, we conducted an evaluation study involving clinical experts. Grounding synthetic data generation in real-world clinical perspectives is crucial for responsible development in AI mental health.
Seven therapists with diverse professional backgrounds (clinical practice, research, education) and extensive experience (6 to 30 years) across various settings (outpatient clinics, hospitals, VA/military, academia) evaluated two full synthetic PE therapy session transcripts. They assessed four key dimensions: content depth, perceived value, session appropriateness, and patient engagement, providing both quantitative ratings and qualitative feedback (See Figure 5).
The evaluation yielded valuable, multifaceted feedback. Experts consistently recognized the dataset’s strengths in capturing detailed patient narratives, with ratings ranging from “Somewhat detailed” (n=3) to “Very detailed” (n=4). Furthermore, the simulated patient’s engagement was rated positively, ranging from “Moderately” engaged (n=4) to “Extremely” engaged (n=3), indicating the model’s success in generating realistic and emotionally resonant patient responses crucial for PE simulations. Identified strengths included vivid trauma descriptions, significant emotional depth, and the portrayal of realistic novice therapist approaches – elements particularly useful for training applications. The feedback also highlighted areas reflecting the inherent challenges of synthetically generating complex therapeutic interactions.
Ratings for perceived value (ranging from “Not valuable” (n=2) to “Valuable” (n=1)) and appropriateness (from “Not appropriate” (n=3) to “Somewhat appropriate” (n=4)) varied, underscoring the difficulty in fully replicating nuanced clinical judgment and meeting all expert expectations synthetically. Similarly, the AI therapist’s perceived skill level (rated between “Novice” (n=2) and “Competent” (n=3)) suggests the current simulation captures less experienced therapist behavior more readily than expert-level interactions.
These critical assessments are invaluable, providing specific, actionable directions for future refinement. Key areas identified for improvement include enhancing conversational flow, reducing therapist interventions and repetitive prompts (e.g., “take a deep breath”), integrating more adaptive AI responses, and further increasing dialogue authenticity. This expert feedback highlights essential targets for advancing the sophistication of generative models in this sensitive domain.
Overall, the evaluation confirms that the LLM-generated transcripts effectively capture core elements of PE therapy sessions, particularly detailed patient narratives and recognizable therapeutic techniques. The diverse evaluator backgrounds provided robust, grounded feedback. While acknowledging the areas identified for continued refinement, this expert validation underscores the dataset’s immediate utility as a valuable resource for developing and testing AI models, studying specific therapeutic phenomena (like emotional arcs or novice therapist patterns), and providing a rich, privacy-preserving alternative to difficult-to-obtain real-world data. This rigorous evaluation process itself demonstrates a commitment to clinical fidelity and transparency in developing AI tools for mental health.
5 Benchmark Setup and Evaluation
This section introduces and validates an emotional trajectory benchmark designed to evaluate AI models’ capabilities in interpreting simulated Prolonged Exposure (PE) therapy conversations.
5.1 Rationale for a PE Therapy Benchmark
Developing a standardized benchmark is crucial for the responsible advancement and reliable comparison of AI models in trauma-focused care. PE therapy relies on carefully tracking and processing patient distress during imaginal exposure. AI tools intended to support PE delivery or therapist training must demonstrate fidelity to these core therapeutic dynamics. A consistent benchmark ensures potential AI applications align with clinical needs and therapeutic principles, facilitates reproducible research, and tracks progress in developing sophisticated AI for mental health support.
5.2 Benchmark Design
Our benchmark assesses emotional intensity across six conversational segments derived from each simulated therapy session, corresponding to different stages within PE therapy:
  1. 1.
    Orientation to Imaginal Exposure
  2. 2.
    Imaginal Exposure Duration
  3. 3.
    Monitoring SUDS Ratings
  4. 4.
    Reinforcing Comments
  5. 5.
    Eliciting Thoughts and Feelings (Processing Step 1)
  6. 6.
    Processing the Imaginal (Processing Step 2)
The expected emotional trajectory follows a recognizable pattern (Cowden Hindash et al., 2020; Bluett et al., 2014; Center for the Treatment and Study of Anxiety, 2023): initial anxiety during orientation, peaking distress during imaginal exposure, followed by gradual reduction through reinforcement and processing (see Figure 6). SUDS (Subjective Units of Distress Scale) are used to measure self-reported distress during exposure.
Although "Monitoring SUDS Ratings" is a distinct segment in our dataset, it occurs concurrently with the "Imaginal Exposure" phase in the PE protocol. Therefore, their emotional intensity ratings are expected to align closely, reflecting shared peak distress dynamics.
5.3 Evaluation Metrics
To quantify how well a model’s predicted emotional trajectory aligns with a baseline, we use three core metrics:
  • Sequence Similarity (Pearson Correlation): Measures the linear correlation between the model’s predicted emotional intensity sequence and the baseline sequence across the six phases. It assesses alignment in both magnitude and relative ranking of distress levels.
  • Pattern Accuracy (Dynamic Time Warping - DTW): Measures the similarity between the two sequences while allowing for non-linear warping in the time dimension. It captures how well the model follows the expected phase-wise progression, even with slight timing shifts. Lower DTW distances indicate better pattern accuracy.
  • Phase Consistency (Root Mean Squared Error - RMSE): Measures the average magnitude of error between the model’s predicted intensity score and the baseline score at each specific therapy phase. Lower RMSE indicates better phase-specific accuracy.
5.4 Baseline Establishment and Validation
To establish a reference for comparison, we generated a Baseline Trajectory using zero-shot emotional intensity ratings from Claude Sonnet 3.5. Models assign scores on a 1 (calm) to 10 (extreme distress) scale for each of the six conversation segments per client profile. Claude Sonnet 3.5 was chosen for its strong zero-shot natural language understanding capabilities, suggesting suitability for interpreting emotional nuances without task-specific training. As shown in Figure 6, the resulting Baseline Trajectory aligns with established PE therapy patterns: peak distress occurs during the Imaginal Exposure and concurrent Monitoring SUDS Ratings phases, with gradual reductions during subsequent Reinforcing, Eliciting Thoughts/Feelings, and Processing phases. This correlates with the human evaluation and confirms the baseline’s suitability for evaluating future models.
5.5 Benchmark Evaluation Methodology
We evaluated AI model performance by comparing their predicted emotional trajectories against the Claude Baseline Trajectory using the defined metrics (Pearson, DTW, RMSE), averaged across all 500 conversation profiles.
To compare models relative to the baseline, we compute performance ratios for each metric (M = model, C = Claude baseline):
  • Pearson (higher is better): $\text{R}_{\text{Corr}}=\frac{\text{M}_{\text{Pearson}}}{\text{C}_{\text{Pearson}}}$
  • DTW (lower is better): $\text{R}_{\text{DTW}}=\frac{\text{C}_{\text{DTW}}}{\text{M}_{\text{DTW}}}$
  • RMSE (lower is better): $\text{R}_{\text{RMSE}}=\frac{\text{C}_{\text{RMSE}}}{\text{M}_{\text{RMSE}}}$
We selected a range of comparison models representing different sizes, architectures, and training methodologies (Mistral Large, Amazon Nova Pro, Llama3 70B/8B Instruct, Llama 3.1 70B/8B Instruct, Mistral 7B Instruct, Mistral Small) to test the benchmark’s ability to differentiate capabilities.
5.6 Results and Discussion
The evaluation results, summarized in Table 1, demonstrate the benchmark’s ability to quantitatively differentiate model performance in assessing emotional trajectories.
To provide a single summary measure, we developed an Absolute Composite Score ($S_{\text{abs}}$). This score combines the normalized and direction-aligned values of the average Pearson correlation ($\uparrow$), DTW distance ($\downarrow$), and RMSE ($\downarrow$) into a single value between 0 and 1 (higher is better). It reflects overall performance relative to theoretical bounds, addressing the challenge of comparing metrics with different scales and optimal directions.
The calculation involved three steps:
  1. 1.
    Normalization to Fixed Bounds (0–1 Scale): Each metric’s average value ($\overline{P}$, $\overline{D}$, $\overline{R}$) was normalized to a scale from 0 (worst bound) to 1 (best bound) using fixed theoretical bounds based on the metric’s properties, resulting in $P_{\text{norm}}$, $D_{\text{norm}}$, and $R_{\text{norm}}$. Values were clipped to $[0,1]$.
    • Pearson Correlation ($P_{\text{norm}}$): Normalized using bounds $[0,1]$, as meaningful correlations range from 0 (no correlation) to 1 (perfect correlation).
      $P_{\text{norm}}=\frac{\overline{P}-0}{1-0}$
    • RMSE ($R_{\text{norm}}$): Normalized using bounds $[0,9.0]$, where 0 represents perfect agreement with the baseline, and 9.0 is the maximum possible RMSE given the 1–10 emotional score range (i.e., $|1-10|=9$).
      $R_{\text{norm}}=\frac{\overline{R}-0}{9.0}$
    • DTW ($D_{\text{norm}}$): Normalized using bounds $[0,5.0]$, where 0 represents identical sequences. The upper bound of 5.0 was selected pragmatically to moderately exceed the maximum observed average $\overline{D}\approx 3.3$, providing sensitivity while ensuring stability against potential outliers.
      $D_{\text{norm}}=\frac{\overline{D}-0}{5.0}$
  2. 2.
    Direction Alignment (Higher = Better): Lower-is-better metrics (RMSE, DTW) were inverted ($X_{\text{aligned}}=1-X_{\text{norm}}$) so that higher values always indicate better performance. Pearson already aligned ($P_{\text{aligned}}=P_{\text{norm}}$). The equations are: $R_{\text{aligned}}=1-R_{\text{norm}}$; $D_{\text{aligned}}=1-D_{\text{norm}}$; $P_{\text{aligned}}=P_{\text{norm}}$
  3. 3.
    Combination (Simple Average): The final score $S_{\text{abs}}$ is the simple average of the aligned, normalized scores, providing a balanced overall measure.
    $S_{\text{abs}}=\frac{P_{\text{aligned}}+R_{\text{aligned}}+D_{\text{aligned}}}{3}$
The resulting $S_{\text{abs}}$ ranges from 0 to 1, indicating a model’s overall alignment (1.0 = ideal) with the baseline trajectory across the three metrics.
Table 1: Benchmark Comparison Results Against Claude Sonnet 3.5 Baseline
Model N* Pearson $\uparrow$ (Avg ± S.D) DTW $\downarrow$ (Avg ± S.D) RMSE $\downarrow$ (Avg ± S.D) $S_{\text{abs}}$uparrow
Mistral Large 500 0.80 ± 0.14 2.38 ± 0.69 1.07 ± 0.33 0.74
Amazon Nova Pro 500 0.74 ± 0.16 2.63 ± 0.73 1.24 ± 0.35 0.69
Llama 3 70B Instruct 489 0.73 ± 0.16 2.61 ± 0.75 1.28 ± 0.36 0.69
Llama 3.1 70B Instruct 500 0.70 ± 0.17 2.80 ± 0.73 1.29 ± 0.35 0.67
Llama 3 8B Instruct 489 0.64 ± 0.23 3.24 ± 0.84 1.61 ± 0.43 0.61
Llama 3.1 8B Instruct 500 0.63 ± 0.23 2.91 ± 0.70 1.44 ± 0.37 0.63
Mistral 7B Instruct 500 0.62 ± 0.21 2.88 ± 0.75 1.49 ± 0.38 0.63
Mistral Small 500 0.61 ± 0.20 3.30 ± 0.94 1.70 ± 0.42 0.59
*N=489 for original Llama 3 v1 models due to limited 8k context window limit exceeded by some samples.
As shown in Table 1, Mistral Large exhibited the strongest alignment ($S_{\text{abs}}=0.74$) with the baseline, achieving the highest Pearson correlation (0.80) and lowest DTW (2.38) and RMSE (1.07), with the lowest standard deviations indicating high consistency. Amazon Nova Pro ($S_{\text{abs}}=0.69$) and Llama 3 70B Instruct ($S_{\text{abs}}=0.69$) performed second best.
Interestingly, Mistral Small ($S_{\text{abs}}=0.59$) showed the weakest alignment, performing worse than the smaller Mistral 7B Instruct model ($S_{\text{abs}}=0.63$). This finding highlights the importance of instruction-following capabilities for the benchmark. Mistral 7B Instruct is specifically tuned for following commands, crucial for adhering to the benchmark’s requirements (e.g., correct scoring scale and format). Mistral Small’s observed difficulties generating correctly formatted responses support the hypothesis that its weaker performance stems from poorer instruction adherence or task suitability rather than the model size. Overall, these comparisons validate the benchmark’s effectiveness in providing quantitative measures of alignment and differentiating between models with varying capabilities for tracking emotional intensity in simulated PE therapy conversations.
6 Future Directions
Future research should aim to extend this benchmark by incorporating other relevant attributes, including granular emotion detection (e.g., distinguishing fear vs. anger), and demographic-based evaluation to assess fairness across different client populations. These extensions are crucial for developing more robust, ethically responsible, and widely applicable systems.
Building on this work, future research should leverage synthetic therapy session conversations to augment real-world datasets, enhancing AI model training with a focus on clinical fidelity. This dataset can serve as a novel tool for therapist education. For example, simulated patient profiles can be used in role-playing, case studies, and interactive simulations, which can lead to more effective therapist training. This dataset can also help develop supportive technologies for patients. For example, this dataset can potentially be extended to facilitate supportive chatbot development.
Chatbots trained on nuanced emotional and therapeutic data can expand access to mental health support, offering anonymous, always-available assistance, especially for those facing barriers like cost, location, or stigma. While not a replacement for human therapists, they can serve as a first point of contact, support mild to moderate concerns, and detect crises, directing users to emergency services.
Future research in this domain can use the dataset to improve model robustness and generalizability. Synthetic datasets can enable researchers to develop and refine algorithms for predicting mental health risks and customizing interventions without compromising privacy.
7 Data Availability
The dataset and code will be made publicly available upon acceptance.
8 Ethical Considerations
This dataset is entirely synthetic and does not involve any real individuals or their personal experiences. While it captures potential patterns within therapy conversations, it should not be interpreted as reflecting any specific population or lived experience.
Our work simulates Prolonged Exposure (PE) therapy for PTSD—a treatment involving sensitive and traumatic content. We recognize the ethical implications related to representation, misuse, and potential harm.
To address these concerns: (1) no real patient data was used—the synthetic dialogues were generated using language models prompted with protocol-adherent scenarios from publicly available PE manuals; (2) we avoid simulating graphic trauma details, focusing on structural and therapeutic fidelity; and (3) any real transcripts used for comparison were collected under IRB-approved protocols with participant consent and de-identified prior to analysis. The dataset release will include usage guidelines and license restrictions to discourage misuse in non-clinical or consumer-facing applications. Our intent is to support clinical NLP research in a privacy-preserving, scalable way—not to replace trained therapists.
We emphasize the need for continued community discussion on the ethical use of synthetic therapy data. Our work does not validate the clinical effectiveness of PE or its synthetic representation. Rather, it investigates whether LLMs can replicate the structure of therapy dialogues for research and training purposes—not simulate real trauma responses or therapeutic outcomes.
9 Limitations
Predefined categories and distributions may not capture the full complexity of real-world therapeutic interactions. The probabilistic nature of the generation process introduces a degree of randomness that may affect the reproducibility of specific analyses. Furthermore, the dataset is limited to the variables and relationships explicitly defined in the generation script.
10 Acknowledgement
This work is supported by the National Science Foundation (NSF) under Grant No. 2326144. Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.
Appendix A Prompt Templates
We include the full prompts used to generate the synthetic therapy transcripts across different stages of Prolonged Exposure (PE) therapy. Each prompt was designed to elicit realistic, structured, and therapeutically valid conversations from a Large Language Model.
A.1 Prompt P5: Orientation to Imaginal Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue where a therapist orients the client to the imaginal exposure planned for a Prolonged Exposure (PE) therapy session.
Key Features:
- The therapist explains the purpose and rationale behind imaginal exposure in a clear and empathetic manner.
- The therapist addresses the clients concerns (if any), hesitations (if any), or questions about the exercise.
- Include the therapist setting expectations for the session, including what the client might feel and how they will be supported throughout.
- The dialogue should include the clients responses, such as questions, emotional reactions, or expressions of understanding.
- The therapist reinforces the importance of the exercise in addressing PTSD symptoms and validates the clients courage in participating.
- Ensure the conversation flows naturally, with pauses, realistic emotional exchanges, and detailed descriptions of both the therapists and clients perspectives.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.2 Prompt P6: Monitoring SUDS Ratings
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue from a Prolonged Exposure (PE) therapy session focusing on the therapist monitoring Subjective Units of Distress (SUDS) ratings during an imaginal exposure exercise.
Key Features:
- The therapist asks the client to provide SUDS ratings approximately every 5 minutes.
- The therapist responds empathetically to changes in the clients ratings, showing curiosity and support.
- Include the client describing their emotions, physical sensations, and distress levels in response to the memory.
- The therapist normalizes the clients experience and encourages them to stay engaged, even as distress levels fluctuate.
- Ensure the conversation feels natural, with pauses, filler words, and realistic emotional exchanges.
- Include vivid descriptions of the clients reactions and the therapists responses.
- The session should convey a balance of emotional support and professional guidance.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.3 Prompt P7: Reinforcing During Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue between a therapist and a client during a Prolonged Exposure (PE) therapy session, focusing on the therapist providing reinforcing comments during imaginal exposure.
Key Features:
- The therapist uses appropriate reinforcing comments, such as "Youre doing great," "Stay with it," or "Its okay to feel this way - youre safe here."
- Include moments where the client hesitates, experiences emotional reactions, or struggles, with the therapist providing timely and empathetic reinforcement.
- Reinforce the clients ability to handle difficult emotions and encourage them to stay present in the memory.
- Ensure that reinforcement is balanced with professional boundaries to make the client feel supported but not pressured.
- The dialogue should feel realistic and empathetic, with the therapist validating the clients efforts and guiding them through moments of distress.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. If you need to end, dont end it abruptly. Dont give any text apart from the therapist or client.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.4 Prompt P8: Eliciting Thoughts and Feelings
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue from a Prolonged Exposure (PE) therapy session where the therapist elicits the clients thoughts and feelings during and after an imaginal exposure exercise.
Key Features:
- The therapist uses open-ended questions to encourage the client to reflect on their thoughts and feelings, such as "Whats coming up for you now?" or "What are you feeling in this moment?"
- Include the clients detailed reflections on their emotions, physical sensations, and thoughts in response to the memory.
- The therapist connects the clients thoughts and feelings to their broader trauma experience and recovery journey.
- The therapist provides empathetic and insightful responses to encourage deeper exploration.
- Ensure the dialogue feels natural, with pauses and filler words, and conveys the therapists empathy and professionalism.
- Include vivid descriptions of the clients emotional and cognitive responses to the memory.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. If you need to end, dont end it abruptly. Dont give any text apart from the therapist or client.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.5 Prompt P10: Full Imaginal Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a vivid and detailed imaginal exposure dialogue between a therapist and a client in a Prolonged Exposure (PE) therapy session.
Key Features:
- The client expresses their emotional state in their own words, which may include nervousness, excitement, hesitation, or determination. Avoid repetitive patterns like always starting with "Im nervous."
- The therapist monitors the clients distress and provides supportive interventions (e.g., grounding techniques, encouraging present-tense narration).
- Include moments where the client struggles emotionally or physically, and the therapist responds with empathy and encouragement to keep them engaged.
- Highlight the therapists use of SUDS monitoring and reinforcing comments to guide the client through the exercise.
- The transcript should include natural pauses, filler words, and a balance between vivid client narration and therapeutic intervention.
- Focus on maintaining authenticity and depth throughout.
- Ensure the duration of the dialogue realistically represents the imaginal exposure process and dont stop in between to ask if you need to continue. Just keep going. (about 30-45 minutes).
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.6 Prompt P11: Processing the Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue where a therapist processes the imaginal exposure with the client in a Prolonged Exposure (PE) therapy session.
Key Features:
- The therapist guides the client in reflecting on their emotional and cognitive responses to the imaginal exposure.
- Include open-ended questions from the therapist, such as, "What stood out to you about that experience?" or "How did it feel to go through that memory today?"
- The therapist helps the client connect their reactions during the imaginal to their broader PTSD symptoms and recovery goals.
- Include moments where the client shares their insights, struggles, or progress, and the therapist validates their effort and progress.
- Highlight any specific strategies or learnings that come out of the discussion, and ensure the therapist encourages the clients continued engagement in the therapy process.
- Ensure the conversation feels empathetic, insightful, and natural, with pauses, filler words, and realistic emotional exchanges.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist: