### v3

# Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations

Abstract
The advancement of AI systems for mental health support is hindered by limited access to therapeutic conversation data, particularly for trauma treatment. We present Thousand Voices of Trauma, a synthetic benchmark dataset of 3,000 therapy conversations based on Prolonged Exposure therapy protocols for Post-traumatic Stress Disorder (PTSD). The dataset comprises 500 unique cases, each explored through six conversational perspectives that mirror the progression of therapy from initial anxiety to peak distress to emotional processing. We incorporated diverse demographic profiles (ages 18-80, M=49.3, 49.4% male, 44.4% female, 6.2% non-binary), 20 trauma types, and 10 trauma-related behaviors using deterministic and probabilistic generation methods. Analysis reveals realistic distributions of trauma types (witnessing violence 10.6%, bullying 10.2%) and symptoms (nightmares 23.4%, substance abuse 20.8%). Clinical experts validated the dataset’s therapeutic fidelity, highlighting its emotional depth while suggesting refinements for greater authenticity. We also developed an emotional trajectory benchmark with standardized metrics for evaluating model responses. This privacy-preserving dataset addresses critical gaps in trauma-focused mental health data, offering a valuable resource for advancing both patient-facing applications and clinician training tools.
Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations
Suhas BN1  Dominik Mattioli1  Saeed Abdullah1 Rosa I. Arriaga2  Chris W. Wiese3  Andrew M. Sherrill4 1College of Information Sciences and Technology, Penn State University, USA 2School of Interactive Computing, Georgia Tech, USA 3School of Psychology, Georgia Tech, USA 4Department of Psychiatry and Behavioral Sciences, Emory University, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
1 Introduction
The intersection of mental health care and artificial intelligence presents unprecedented opportunities alongside significant challenges. AI system development faces particular obstacles in trauma-focused therapy due to the sensitive nature of patient experiences and strict privacy regulations, which make the collection of real-world data extremely challenging (Mathur et al., 2022).
Moreover, existing datasets frequently lack the diversity and clinical depth needed to train robust AI systems capable of serving diverse populations effectively Marr (2023). Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD) (Foa et al., 2007)—offers a structured therapeutic approach that could especially benefit from AI support.
However, current mental health conversation datasets are often too small (Zantvoort et al., 2024), lack demographic diversity (Sadeh-Sharvit, 2024), and do not capture the nuanced progression of trauma-focused therapy sessions (Mathur et al., 2022).
To address these gaps, we introduce Thousand Voices of Trauma, a synthetic benchmark dataset comprising 500 clinical sessions. Each session is divided into six distinct therapeutic phases aligned with core evaluation criteria of PE therapy. Each phase contains multiple therapist-client exchanges, which can be analyzed either independently or as part of the complete session flow, resulting in 3000 distinct clinical conversations. These phases, as defined by Foa et al. (2007), correspond to specific stages within the PE therapy session: a) Orientation to Imaginal Exposure, b) Imaginal Exposure Duration, c) Monitoring SUDS Ratings, d) Reinforcing Comments, e) Eliciting Thoughts and Feelings, and f) Processing the Imaginal.
These phases mirror the typical session progression—from initial anxiety, through peak distress during imaginal exposure, to the gradual reduction of distress through reinforcement and processing. Its diversity encompasses a wide range of demographic profiles, trauma types, and associated behaviors, designed to reflect varied real-world clinical presentations. This structured, diverse dataset offers scalable opportunities for AI systems to assist mental health professionals in trauma-focused therapy. By emphasizing diverse populations, Thousand Voices of Trauma represents a meaningful step toward more effective, personalized, and ethically guided mental health care.
This dataset also addresses other real-world limitations that often hinder mental health research and safe AI model development. For example, while privacy concerns often restrict data access, synthetic data can circumvent typical ethical and legal barriers. It also overcomes other common issues with real-world data, such as incompleteness, inconsistency, and small sample sizes—especially among minority groups.
By balancing representation across diverse populations, trauma types, and racial or ethnic minorities, the dataset helps mitigate inherent biases. For instance, the NIMH reported in 2021 that 14.5 million U.S. adults (5.7%) experienced severe major depressive episodes, with higher rates among females (10.3%) than males (6.2%) and the highest prevalence among those aged 18–25 (18.6%). Synthetic data can compensate for such imbalances, enhancing model training and analysis.
1.1 Key Contributions
  1. 1.
    Scale and Diversity: Our dataset is one of the largest collections of synthetic trauma-focused therapy conversations, encompassing varied demographic profiles across age, gender, ethnicity, and cultural backgrounds. The use of synthetic data addresses ethical concerns related to patient privacy and consent, while its diversity supports the development of more inclusive and culturally competent AI systems.
  2. 2.
    Clinical Depth: Each conversation is grounded in evidence-based PE therapy principles and incorporates 20 distinct trauma types, 10 trauma-related behaviors, and 5 co-occurring conditions. This clinical specificity renders the dataset particularly valuable for specialized therapeutic applications, including clinician training.
  3. 3.
    Structured Evaluation Framework: The six-conversation structure per session enables a comprehensive examination of different interaction trajectories, spanning from the initial assessment to trauma processing and progress evaluation.
  4. 4.
    Baseline: We establish a baseline dataset for training AI systems in trauma-focused therapy, facilitating standardized comparisons across different approaches.
The remainder of this paper is structured as follows: Section 2 reviews related work in mental health. Section 3 describes the datasets and details the synthetic data generation methodology. Section 4 presents the expert evaluation. Section 5 introduces the benchmark setup and evaluation method. Sections 69 discuss future directions, data availability, ethical considerations, and limitations.
2 Related Work
Prolonged Exposure (PE) therapy, an evidence-based treatment for PTSD, relies on structured exposure to trauma-related narratives (Watkins et al., 2018). However, there is a lack of trained professionals who can provide PE therapy (Rauch et al., 2017, 2023). As a result, there is an urgent need for AI applications to support PE therapy delivery and training. This underscores the need for clinically valid and diverse datasets for AI development and evaluation. Large-scale language models (LLMs), like the GPT series, have shown potential in generating synthetic datasets that mimic human-like text, addressing challenges such as data scarcity and privacy concerns (Hämäläinen et al., 2023; Giuffrè and Shung, 2023; Zhu, 2024; Li et al., 2024). However, for applications like PE therapy—which require alignment with trauma-focused frameworks, diverse demographic representation, and strict ethical safeguards—current research still lacks tailored solutions.
Synthetic datasets show promise in mental health applications, with studies exploring LLM-based data generation to address data scarcity. Wu et al. (Wu et al., 2023, 2024) introduced zero-shot and few-shot learning frameworks to augment PTSD diagnostic datasets, producing synthetic transcripts that outperform baselines. The latter work used role-prompting and structured prompts to create realistic synthetic clinical interviews.
However, these approaches have not been specifically adapted to the structured narratives essential for PE therapy. Efforts to enhance demographic diversity in synthetic datasets are growing. Mori et al. (2024) and Lozoya et al. (2023) examined how synthetic data reflects demographic variation, highlighting biases in LLM outputs, especially regarding race and gender, and stressing fairness in mental health datasets. Techniques like patient vignette simulation (Denecke, 2024) and adaptive prompts for non-English contexts (e.g., SAPE for Spanish (Lozoya et al., 2024)) show early progress toward inclusivity. However, trauma-type diversity and PE-specific scenarios remain unexplored. Additionally, Chen et al. (2024) underscored the need for systematic benchmarking using metrics like F1-score, AUC, and balanced accuracy, but these tools have not yet been applied to datasets focused on diverse trauma types or PE therapy.
Privacy and ethics are central to synthetic dataset generation, with studies focusing on privacy-preserving methods. Recent works (Chuang et al., 2024a; Meoni et al., 2024; Suhas et al., 2023; BN and Abdullah, 2022; Chuang et al., 2024b) highlight privacy-preserving machine learning, protected health information (PHI) exclusion, and semantic filtering to maintain privacy compliance while preserving data utility. While these methods offer strong safeguards for general clinical use, ethical risks—such as generating harmful trauma narratives or victim-blaming—remain underexplored, especially in sensitive contexts like trauma-focused therapies such as PE.
While synthetic dataset generation using LLMs (Wu et al., 2023, 2024; Xu et al., 2024) has advanced in addressing general clinical challenges, gaps in the literature remain for PE therapy. These include the lack of trauma-type diversity, limited demographic inclusivity evaluation, insufficient alignment with frameworks like DSM-5 PTSD criteria, and underdeveloped ethical safeguards specific to trauma-focused contexts. Synthetic datasets can also help mitigate representation biases in AI models for mental healthcare delivery.
For example, American Psychiatric Association (2013) reported that non-Hispanic White adults (25.0%) were more likely to receive mental health services than non-Hispanic Black (18.3%), Hispanic (17.3%), and Asian (13.9%) adults. Including such underrepresented groups in synthetic datasets might partially address the training data gap. This paper seeks to bridge these gaps by exploring the existing knowledge base and identifying pathways to tailor synthetic data generation for PE therapy applications.
Figure 1: The figure presents the demographic distribution of generated participant profiles across four categories: gender, age, ethnicity, and relationship status. The gender distribution shows a nearly balanced representation of males (247) and females (222), with a smaller group identifying as non-binary (31) mimicking real world gender distributions U.S. Census Bureau (2025). The Age Distribution spans from under 10 to over 90 years, with most participants between 30 and 70 years. The Ethnicity Distribution highlights diverse backgrounds, with the largest groups being Latin American (70), North American (67), South East Asian (64), South Asian (63), and European (63), followed by Oceania (62), Middle Eastern (59), and African (52) participants. The Relationship Status chart reveals that most participants are married (180) or single (149), while smaller numbers are in relationships (70), divorced (51), or widowed (50).
Figure 2: The figure presents the distribution of exhibited behaviors and types of trauma among participants. The Exhibited Behaviors chart shows that nightmares (117 participants) are the most common symptom, followed by substance abuse (104), compulsive behaviors (101), aggression (100), and hypervigilance (99). Other frequently reported behaviors include self-blame (97), withdrawal (95), flashbacks (95), avoidance (91), and dissociation (83). The Types of Trauma chart highlights that witnessing violence (53 participants) and bullying (51) are the most frequently reported traumatic experiences. Other significant trauma types include neglect (45), medical trauma (43), childhood trauma (42), and the loss of a loved one (39). Natural disasters, sexual abuse, physical abuse, and emotional abuse were each reported by 39 to 38 participants, while combat or war experiences and accidents were each cited by 36 participants.
3 Dataset
3.1 Dataset: Simulated Therapy Session Profiles
To systematically examine trauma narratives and therapeutic dynamics, we constructed a synthetic dataset comprising 3,000 simulated therapy conversations. These dialogues were generated using Claude Sonnet 3.5 (Anthropic, 2024), selected as the foundational model for the Thousand Voices of Trauma dataset based on its demonstrated ability to produce coherent, contextually grounded, and richly detailed dialogue during preliminary assessments. The rationale for prioritizing Claude Sonnet 3.5 was centered on its superior generative capacity, which allowed the efficient production of high-quality conversations that aligned closely with the principles of prolonged exposure therapy (PE). This choice allowed us to pay more attention to the methodological challenges of scenario construction, prompt engineering, and the integration of PTSD-specific symptomatology and therapeutic constructs.
Although comparative evaluation of other large language models remains a compelling direction for future investigation, the initial deployment of a robust, high-performing model was critical to establishing a reliable and practically useful foundation for trauma-focused research. This decision facilitated the immediate applicability of the dataset for downstream tasks, including the modeling of trauma-informed therapeutic interactions and the development of evaluation frameworks for clinical AI tools.
Each simulated profile integrates structured metadata, including client demographics (e.g., age, gender, living situation), therapist attributes, and session-level variables such as trauma type, therapeutic phase, and discussion topics. To ensure diversity and ecological validity, both deterministic (rule-based) and probabilistic (sampling-based) generation methods were used in constructing the scenarios.
To illustrate model-specific stylistic and structural variations, we provide a comparative set of examples generated by multiple state-of-the-art Frontier Models. Full prompt templates and representative conversations for each phase of the therapeutic protocol are included in Appendix A and B.
3.2 Client Profile Generation
Client profiles included age, gender, relationship status, occupation, living situation, and ethnicity. Ages ranged from 18 to 80, divided into six groups: 18-30, 31-40, 41-50, 51-60, 61-70, 71-80. We assigned gender using weighted probabilities: 50% male, 49% female, and 1% non-binary U.S. Census Bureau (2025). Relationship status, occupation, and living situation were age-specific—for example, clients aged 20-30 were more likely to be “Single,” “Student,” and “With parents,” while those 60-70 were often “Widowed,” “Retired,” and “Alone.” A validation function ensured logical consistency. We randomly assigned ethnicity from eight global regions: South Asian, Middle Eastern, African, North American, Oceanian, European, South East Asian, and Latin American. We assigned co-occurring conditions with weighted probabilities Bilevicius et al. (2020); Jennifer et al. (2024); Hagiwara et al. (2022): None (25%), Anxiety (25%), Depression (30%), Substance Use Disorder (10%), and Chronic Pain (10%). We also assigned clients 1–3 trauma-related behaviors from ten options, including avoidance, hypervigilance, flashbacks, nightmares, self-blame, substance abuse, aggression, withdrawal, dissociation, and compulsive behaviors. The options represent a range of cognitive, emotional, and behavioral responses typically associated with trauma, which aligns with trauma-informed care principles (SAMHSA, 2015, 2014).
3.3 Therapist Profile Generation
We generated therapist profiles with ages ranging from 25 to 65, divided into four age groups: 25-34, 35-44, 45-54, and 55-65. We assigned therapist gender using the same weighted probabilities as client gender.
3.4 Therapy Context Generation
To create varied therapeutic contexts, we defined both a broad trauma type representing the general category of traumatic experience, and a more specific session topic providing the focus for the simulated interaction. We generated the therapy context by randomly selecting a trauma type from twelve possibilities (SAMHSA, 2015, 2014): physical abuse, emotional abuse, sexual abuse, neglect, natural disasters, accidents, combat or war experiences, loss of a loved one, witnessing violence, bullying, childhood trauma, and medical trauma. A session topic was also randomly selected from twenty possibilities, such as car accident trauma, workplace trauma, domestic violence, natural disaster experience, military combat experience, loss of a loved one, severe illness, divorce or relationship breakup, witnessing violence, racial trauma, and refugee experiences. This independent random selection of trauma type and session topic allows for a wide range of generated scenarios. While this methodology can result in pairings without an immediate thematic link (e.g., ‘natural disaster’ type with ‘workplace trauma’ topic), such combinations can also reflect the complexity observed in real-world therapy, where session discussions may focus on various life stressors influenced by, or co-existing with, the client’s primary trauma history.
3.5 Session Profile Assembly
Each complete session profile combined a validated client profile, a therapist profile, and the generated therapy context, including the trauma type and session topic.
3.6 Dataset Statistics
The dataset comprises 500 simulated participants (ages 18-80 years, M = 49.3). The gender distribution includes 247 male (49.4%), 222 female (44.4%), and 31 non-binary (6.2%) participants.
The ethnicity distribution is as follows: Latin American (70, 14.0%), North American (67, 13.4%), South East Asian (64, 12.8%), South Asian (63, 12.6%), European (63, 12.6%), Oceania (62, 12.4%), Middle Eastern (59, 11.8%), and African (52, 10.4%). Regarding relationship status, participants were predominantly married (180, 36.0%) or single (149, 29.8%), with others reporting being in a relationship (70, 14.0%), divorced (51, 10.2%), or widowed (50, 10.0%). See Figure 1 & 2 for more details.
Figure 3: The figures illustrate structure and language diversity in synthetic therapist-client dialogues. The Utterance Length Distribution (top) shows clients often speak at length (>50 words), while therapists’ responses are concise, reflecting the client-centered nature of therapy. The Vocabulary Diversity (bottom) reveals clients use ~24,000 unique words, far more than therapists (~5,000), likely due to personal narratives, whereas therapists maintain structured, reflective language.
Figure 4: The figure depicts conversation flow in synthetic dialogues, showing exchange lengths over time across three phases: Setup, Exposure, and Processing. In Setup, lengths remain stable (~40 to 45 words). Exposure sees a steady increase, peaking at ~60 words, indicating deeper engagement. Processing shows fluctuations, reflecting varying reflection and emotional processing. The shaded region represents variability across conversations.
Generated interactions exhibited various trauma-related behaviors (See Figure 2, 3 and 4 respectively), with nightmares being most prevalent (117, 23.4%), followed by substance abuse (104, 20.8%), compulsive behaviors (101, 20.2%), and aggression (100, 20.0%). Other common manifestations included hypervigilance (99, 19.8%), self-blame (97, 19.4%), withdrawal and flashbacks (95 each, 19.0%), avoidance (91, 18.2%), and dissociation (83, 16.6%). The types of trauma are diverse, with witnessing violence being most common (53, 10.6%), followed by bullying (51, 10.2%), neglect (45, 9.0%), medical trauma (43, 8.6%), and childhood trauma (42, 8.4%). Other reported traumas included loss of a loved one, natural disasters, sexual abuse, and physical abuse (39 each, 7.8%), emotional abuse (38, 7.6%), and combat or war experiences and accidents (36 each, 7.2%).
Figure 5: Stacked Bar charts illustrating therapist ratings (N=7) across four key dimensions of the synthetic PE therapy sessions—Content Depth, Perceived Value, Session Appropriateness, and Patient Engagement.
Figure 6: Violin plots illustrating the distribution of emotional intensity scores across six therapy stages scored by Claude Sonnet 3.5. Emotional intensity generally starts at moderate levels during orientation, peaks during exposure and monitoring, and then tapers off through reinforcing, reflection, and processing phases.
4 Expert Evaluation of Synthetic PE Therapy Sessions
To ensure clinical relevance and gain insights into the utility and limitations of our synthetic dataset, we conducted an evaluation study involving clinical experts. Grounding synthetic data generation in real-world clinical perspectives is crucial for responsible development in AI mental health.
Seven therapists with diverse professional backgrounds (clinical practice, research, education) and extensive experience (6 to 30 years) across various settings (outpatient clinics, hospitals, VA/military, academia) evaluated two full synthetic PE therapy session transcripts. They assessed four key dimensions: content depth, perceived value, session appropriateness, and patient engagement, providing both quantitative ratings and qualitative feedback (See Figure 5).
The evaluation yielded valuable, multifaceted feedback. Experts consistently recognized the dataset’s strengths in capturing detailed patient narratives, with ratings ranging from “Somewhat detailed” (n=3) to “Very detailed” (n=4). Furthermore, the simulated patient’s engagement was rated positively, ranging from “Moderately” engaged (n=4) to “Extremely” engaged (n=3), indicating the model’s success in generating realistic and emotionally resonant patient responses crucial for PE simulations. Identified strengths included vivid trauma descriptions, significant emotional depth, and the portrayal of realistic novice therapist approaches – elements particularly useful for training applications. The feedback also highlighted areas reflecting the inherent challenges of synthetically generating complex therapeutic interactions.
Ratings for perceived value (ranging from “Not valuable” (n=2) to “Valuable” (n=1)) and appropriateness (from “Not appropriate” (n=3) to “Somewhat appropriate” (n=4)) varied, underscoring the difficulty in fully replicating nuanced clinical judgment and meeting all expert expectations synthetically. Similarly, the AI therapist’s perceived skill level (rated between “Novice” (n=2) and “Competent” (n=3)) suggests the current simulation captures less experienced therapist behavior more readily than expert-level interactions.
These critical assessments are invaluable, providing specific, actionable directions for future refinement. Key areas identified for improvement include enhancing conversational flow, reducing therapist interventions and repetitive prompts (e.g., “take a deep breath”), integrating more adaptive AI responses, and further increasing dialogue authenticity. This expert feedback highlights essential targets for advancing the sophistication of generative models in this sensitive domain.
Overall, the evaluation confirms that the LLM-generated transcripts effectively capture core elements of PE therapy sessions, particularly detailed patient narratives and recognizable therapeutic techniques. The diverse evaluator backgrounds provided robust, grounded feedback. While acknowledging the areas identified for continued refinement, this expert validation underscores the dataset’s immediate utility as a valuable resource for developing and testing AI models, studying specific therapeutic phenomena (like emotional arcs or novice therapist patterns), and providing a rich, privacy-preserving alternative to difficult-to-obtain real-world data. This rigorous evaluation process itself demonstrates a commitment to clinical fidelity and transparency in developing AI tools for mental health.
5 Benchmark Setup and Evaluation
This section introduces and validates an emotional trajectory benchmark designed to evaluate AI models’ capabilities in interpreting simulated Prolonged Exposure (PE) therapy conversations.
5.1 Rationale for a PE Therapy Benchmark
Developing a standardized benchmark is crucial for the responsible advancement and reliable comparison of AI models in trauma-focused care. PE therapy relies on carefully tracking and processing patient distress during imaginal exposure. AI tools intended to support PE delivery or therapist training must demonstrate fidelity to these core therapeutic dynamics. A consistent benchmark ensures potential AI applications align with clinical needs and therapeutic principles, facilitates reproducible research, and tracks progress in developing sophisticated AI for mental health support.
5.2 Benchmark Design
Our benchmark assesses emotional intensity across six conversational segments derived from each simulated therapy session, corresponding to different stages within PE therapy:
  1. 1.
    Orientation to Imaginal Exposure
  2. 2.
    Imaginal Exposure Duration
  3. 3.
    Monitoring SUDS Ratings
  4. 4.
    Reinforcing Comments
  5. 5.
    Eliciting Thoughts and Feelings (Processing Step 1)
  6. 6.
    Processing the Imaginal (Processing Step 2)
The expected emotional trajectory follows a recognizable pattern (Cowden Hindash et al., 2020; Bluett et al., 2014; Center for the Treatment and Study of Anxiety, 2023): initial anxiety during orientation, peaking distress during imaginal exposure, followed by gradual reduction through reinforcement and processing (see Figure 6). SUDS (Subjective Units of Distress Scale) are used to measure self-reported distress during exposure.
Although "Monitoring SUDS Ratings" is a distinct segment in our dataset, it occurs concurrently with the "Imaginal Exposure" phase in the PE protocol. Therefore, their emotional intensity ratings are expected to align closely, reflecting shared peak distress dynamics.
5.3 Evaluation Metrics
To quantify how well a model’s predicted emotional trajectory aligns with a baseline, we use three core metrics:
  • Sequence Similarity (Pearson Correlation): Measures the linear correlation between the model’s predicted emotional intensity sequence and the baseline sequence across the six phases. It assesses alignment in both magnitude and relative ranking of distress levels.
  • Pattern Accuracy (Dynamic Time Warping - DTW): Measures the similarity between the two sequences while allowing for non-linear warping in the time dimension. It captures how well the model follows the expected phase-wise progression, even with slight timing shifts. Lower DTW distances indicate better pattern accuracy.
  • Phase Consistency (Root Mean Squared Error - RMSE): Measures the average magnitude of error between the model’s predicted intensity score and the baseline score at each specific therapy phase. Lower RMSE indicates better phase-specific accuracy.
5.4 Baseline Establishment and Validation
To establish a reference for comparison, we generated a Baseline Trajectory using zero-shot emotional intensity ratings from Claude Sonnet 3.5. Models assign scores on a 1 (calm) to 10 (extreme distress) scale for each of the six conversation segments per client profile. Claude Sonnet 3.5 was chosen for its strong zero-shot natural language understanding capabilities, suggesting suitability for interpreting emotional nuances without task-specific training. As shown in Figure 6, the resulting Baseline Trajectory aligns with established PE therapy patterns: peak distress occurs during the Imaginal Exposure and concurrent Monitoring SUDS Ratings phases, with gradual reductions during subsequent Reinforcing, Eliciting Thoughts/Feelings, and Processing phases. This correlates with the human evaluation and confirms the baseline’s suitability for evaluating future models.
5.5 Benchmark Evaluation Methodology
We evaluated AI model performance by comparing their predicted emotional trajectories against the Claude Baseline Trajectory using the defined metrics (Pearson, DTW, RMSE), averaged across all 500 conversation profiles.
To compare models relative to the baseline, we compute performance ratios for each metric (M = model, C = Claude baseline):
  • Pearson (higher is better): $\text{R}_{\text{Corr}}=\frac{\text{M}_{\text{Pearson}}}{\text{C}_{\text{Pearson}}}$
  • DTW (lower is better): $\text{R}_{\text{DTW}}=\frac{\text{C}_{\text{DTW}}}{\text{M}_{\text{DTW}}}$
  • RMSE (lower is better): $\text{R}_{\text{RMSE}}=\frac{\text{C}_{\text{RMSE}}}{\text{M}_{\text{RMSE}}}$
We selected a range of comparison models representing different sizes, architectures, and training methodologies (Mistral Large, Amazon Nova Pro, Llama3 70B/8B Instruct, Llama 3.1 70B/8B Instruct, Mistral 7B Instruct, Mistral Small) to test the benchmark’s ability to differentiate capabilities.
5.6 Results and Discussion
The evaluation results, summarized in Table 1, demonstrate the benchmark’s ability to quantitatively differentiate model performance in assessing emotional trajectories.
To provide a single summary measure, we developed an Absolute Composite Score ($S_{\text{abs}}$). This score combines the normalized and direction-aligned values of the average Pearson correlation ($\uparrow$), DTW distance ($\downarrow$), and RMSE ($\downarrow$) into a single value between 0 and 1 (higher is better). It reflects overall performance relative to theoretical bounds, addressing the challenge of comparing metrics with different scales and optimal directions.
The calculation involved three steps:
  1. 1.
    Normalization to Fixed Bounds (0–1 Scale): Each metric’s average value ($\overline{P}$, $\overline{D}$, $\overline{R}$) was normalized to a scale from 0 (worst bound) to 1 (best bound) using fixed theoretical bounds based on the metric’s properties, resulting in $P_{\text{norm}}$, $D_{\text{norm}}$, and $R_{\text{norm}}$. Values were clipped to $[0,1]$.
    • Pearson Correlation ($P_{\text{norm}}$): Normalized using bounds $[0,1]$, as meaningful correlations range from 0 (no correlation) to 1 (perfect correlation).
      $P_{\text{norm}}=\frac{\overline{P}-0}{1-0}$
    • RMSE ($R_{\text{norm}}$): Normalized using bounds $[0,9.0]$, where 0 represents perfect agreement with the baseline, and 9.0 is the maximum possible RMSE given the 1–10 emotional score range (i.e., $|1-10|=9$).
      $R_{\text{norm}}=\frac{\overline{R}-0}{9.0}$
    • DTW ($D_{\text{norm}}$): Normalized using bounds $[0,5.0]$, where 0 represents identical sequences. The upper bound of 5.0 was selected pragmatically to moderately exceed the maximum observed average $\overline{D}\approx 3.3$, providing sensitivity while ensuring stability against potential outliers.
      $D_{\text{norm}}=\frac{\overline{D}-0}{5.0}$
  2. 2.
    Direction Alignment (Higher = Better): Lower-is-better metrics (RMSE, DTW) were inverted ($X_{\text{aligned}}=1-X_{\text{norm}}$) so that higher values always indicate better performance. Pearson already aligned ($P_{\text{aligned}}=P_{\text{norm}}$). The equations are: $R_{\text{aligned}}=1-R_{\text{norm}}$; $D_{\text{aligned}}=1-D_{\text{norm}}$; $P_{\text{aligned}}=P_{\text{norm}}$
  3. 3.
    Combination (Simple Average): The final score $S_{\text{abs}}$ is the simple average of the aligned, normalized scores, providing a balanced overall measure.
    $S_{\text{abs}}=\frac{P_{\text{aligned}}+R_{\text{aligned}}+D_{\text{aligned}}}{3}$
The resulting $S_{\text{abs}}$ ranges from 0 to 1, indicating a model’s overall alignment (1.0 = ideal) with the baseline trajectory across the three metrics.
Table 1: Benchmark Comparison Results Against Claude Sonnet 3.5 Baseline
Model N* Pearson $\uparrow$ (Avg ± S.D) DTW $\downarrow$ (Avg ± S.D) RMSE $\downarrow$ (Avg ± S.D) $S_{\text{abs}}$uparrow
Mistral Large 500 0.80 ± 0.14 2.38 ± 0.69 1.07 ± 0.33 0.74
Amazon Nova Pro 500 0.74 ± 0.16 2.63 ± 0.73 1.24 ± 0.35 0.69
Llama 3 70B Instruct 489 0.73 ± 0.16 2.61 ± 0.75 1.28 ± 0.36 0.69
Llama 3.1 70B Instruct 500 0.70 ± 0.17 2.80 ± 0.73 1.29 ± 0.35 0.67
Llama 3 8B Instruct 489 0.64 ± 0.23 3.24 ± 0.84 1.61 ± 0.43 0.61
Llama 3.1 8B Instruct 500 0.63 ± 0.23 2.91 ± 0.70 1.44 ± 0.37 0.63
Mistral 7B Instruct 500 0.62 ± 0.21 2.88 ± 0.75 1.49 ± 0.38 0.63
Mistral Small 500 0.61 ± 0.20 3.30 ± 0.94 1.70 ± 0.42 0.59
*N=489 for original Llama 3 v1 models due to limited 8k context window limit exceeded by some samples.
As shown in Table 1, Mistral Large exhibited the strongest alignment ($S_{\text{abs}}=0.74$) with the baseline, achieving the highest Pearson correlation (0.80) and lowest DTW (2.38) and RMSE (1.07), with the lowest standard deviations indicating high consistency. Amazon Nova Pro ($S_{\text{abs}}=0.69$) and Llama 3 70B Instruct ($S_{\text{abs}}=0.69$) performed second best.
Interestingly, Mistral Small ($S_{\text{abs}}=0.59$) showed the weakest alignment, performing worse than the smaller Mistral 7B Instruct model ($S_{\text{abs}}=0.63$). This finding highlights the importance of instruction-following capabilities for the benchmark. Mistral 7B Instruct is specifically tuned for following commands, crucial for adhering to the benchmark’s requirements (e.g., correct scoring scale and format). Mistral Small’s observed difficulties generating correctly formatted responses support the hypothesis that its weaker performance stems from poorer instruction adherence or task suitability rather than the model size. Overall, these comparisons validate the benchmark’s effectiveness in providing quantitative measures of alignment and differentiating between models with varying capabilities for tracking emotional intensity in simulated PE therapy conversations.
6 Future Directions
Future research should aim to extend this benchmark by incorporating other relevant attributes, including granular emotion detection (e.g., distinguishing fear vs. anger), and demographic-based evaluation to assess fairness across different client populations. These extensions are crucial for developing more robust, ethically responsible, and widely applicable systems.
Building on this work, future research should leverage synthetic therapy session conversations to augment real-world datasets, enhancing AI model training with a focus on clinical fidelity. This dataset can serve as a novel tool for therapist education. For example, simulated patient profiles can be used in role-playing, case studies, and interactive simulations, which can lead to more effective therapist training. This dataset can also help develop supportive technologies for patients. For example, this dataset can potentially be extended to facilitate supportive chatbot development.
Chatbots trained on nuanced emotional and therapeutic data can expand access to mental health support, offering anonymous, always-available assistance, especially for those facing barriers like cost, location, or stigma. While not a replacement for human therapists, they can serve as a first point of contact, support mild to moderate concerns, and detect crises, directing users to emergency services.
Future research in this domain can use the dataset to improve model robustness and generalizability. Synthetic datasets can enable researchers to develop and refine algorithms for predicting mental health risks and customizing interventions without compromising privacy.
7 Data Availability
The dataset and code will be made publicly available upon acceptance.
8 Ethical Considerations
This dataset is entirely synthetic and does not involve any real individuals or their personal experiences. While it captures potential patterns within therapy conversations, it should not be interpreted as reflecting any specific population or lived experience.
Our work simulates Prolonged Exposure (PE) therapy for PTSD—a treatment involving sensitive and traumatic content. We recognize the ethical implications related to representation, misuse, and potential harm.
To address these concerns: (1) no real patient data was used—the synthetic dialogues were generated using language models prompted with protocol-adherent scenarios from publicly available PE manuals; (2) we avoid simulating graphic trauma details, focusing on structural and therapeutic fidelity; and (3) any real transcripts used for comparison were collected under IRB-approved protocols with participant consent and de-identified prior to analysis. The dataset release will include usage guidelines and license restrictions to discourage misuse in non-clinical or consumer-facing applications. Our intent is to support clinical NLP research in a privacy-preserving, scalable way—not to replace trained therapists.
We emphasize the need for continued community discussion on the ethical use of synthetic therapy data. Our work does not validate the clinical effectiveness of PE or its synthetic representation. Rather, it investigates whether LLMs can replicate the structure of therapy dialogues for research and training purposes—not simulate real trauma responses or therapeutic outcomes.
9 Limitations
Predefined categories and distributions may not capture the full complexity of real-world therapeutic interactions. The probabilistic nature of the generation process introduces a degree of randomness that may affect the reproducibility of specific analyses. Furthermore, the dataset is limited to the variables and relationships explicitly defined in the generation script.
10 Acknowledgement
This work is supported by the National Science Foundation (NSF) under Grant No. 2326144. Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.
Appendix A Prompt Templates
We include the full prompts used to generate the synthetic therapy transcripts across different stages of Prolonged Exposure (PE) therapy. Each prompt was designed to elicit realistic, structured, and therapeutically valid conversations from a Large Language Model.
A.1 Prompt P5: Orientation to Imaginal Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue where a therapist orients the client to the imaginal exposure planned for a Prolonged Exposure (PE) therapy session.
Key Features:
- The therapist explains the purpose and rationale behind imaginal exposure in a clear and empathetic manner.
- The therapist addresses the clients concerns (if any), hesitations (if any), or questions about the exercise.
- Include the therapist setting expectations for the session, including what the client might feel and how they will be supported throughout.
- The dialogue should include the clients responses, such as questions, emotional reactions, or expressions of understanding.
- The therapist reinforces the importance of the exercise in addressing PTSD symptoms and validates the clients courage in participating.
- Ensure the conversation flows naturally, with pauses, realistic emotional exchanges, and detailed descriptions of both the therapists and clients perspectives.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.2 Prompt P6: Monitoring SUDS Ratings
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue from a Prolonged Exposure (PE) therapy session focusing on the therapist monitoring Subjective Units of Distress (SUDS) ratings during an imaginal exposure exercise.
Key Features:
- The therapist asks the client to provide SUDS ratings approximately every 5 minutes.
- The therapist responds empathetically to changes in the clients ratings, showing curiosity and support.
- Include the client describing their emotions, physical sensations, and distress levels in response to the memory.
- The therapist normalizes the clients experience and encourages them to stay engaged, even as distress levels fluctuate.
- Ensure the conversation feels natural, with pauses, filler words, and realistic emotional exchanges.
- Include vivid descriptions of the clients reactions and the therapists responses.
- The session should convey a balance of emotional support and professional guidance.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.3 Prompt P7: Reinforcing During Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue between a therapist and a client during a Prolonged Exposure (PE) therapy session, focusing on the therapist providing reinforcing comments during imaginal exposure.
Key Features:
- The therapist uses appropriate reinforcing comments, such as "Youre doing great," "Stay with it," or "Its okay to feel this way - youre safe here."
- Include moments where the client hesitates, experiences emotional reactions, or struggles, with the therapist providing timely and empathetic reinforcement.
- Reinforce the clients ability to handle difficult emotions and encourage them to stay present in the memory.
- Ensure that reinforcement is balanced with professional boundaries to make the client feel supported but not pressured.
- The dialogue should feel realistic and empathetic, with the therapist validating the clients efforts and guiding them through moments of distress.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. If you need to end, dont end it abruptly. Dont give any text apart from the therapist or client.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.4 Prompt P8: Eliciting Thoughts and Feelings
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue from a Prolonged Exposure (PE) therapy session where the therapist elicits the clients thoughts and feelings during and after an imaginal exposure exercise.
Key Features:
- The therapist uses open-ended questions to encourage the client to reflect on their thoughts and feelings, such as "Whats coming up for you now?" or "What are you feeling in this moment?"
- Include the clients detailed reflections on their emotions, physical sensations, and thoughts in response to the memory.
- The therapist connects the clients thoughts and feelings to their broader trauma experience and recovery journey.
- The therapist provides empathetic and insightful responses to encourage deeper exploration.
- Ensure the dialogue feels natural, with pauses and filler words, and conveys the therapists empathy and professionalism.
- Include vivid descriptions of the clients emotional and cognitive responses to the memory.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. If you need to end, dont end it abruptly. Dont give any text apart from the therapist or client.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.5 Prompt P10: Full Imaginal Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a vivid and detailed imaginal exposure dialogue between a therapist and a client in a Prolonged Exposure (PE) therapy session.
Key Features:
- The client expresses their emotional state in their own words, which may include nervousness, excitement, hesitation, or determination. Avoid repetitive patterns like always starting with "Im nervous."
- The therapist monitors the clients distress and provides supportive interventions (e.g., grounding techniques, encouraging present-tense narration).
- Include moments where the client struggles emotionally or physically, and the therapist responds with empathy and encouragement to keep them engaged.
- Highlight the therapists use of SUDS monitoring and reinforcing comments to guide the client through the exercise.
- The transcript should include natural pauses, filler words, and a balance between vivid client narration and therapeutic intervention.
- Focus on maintaining authenticity and depth throughout.
- Ensure the duration of the dialogue realistically represents the imaginal exposure process and dont stop in between to ask if you need to continue. Just keep going. (about 30-45 minutes).
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
A.6 Prompt P11: Processing the Exposure
You are an expert in medicine and NLP. Generate a clinical transcript for the following profiles:
<profiles>
{profile_info}
</profiles>
Based on these expectations:
Generate a detailed creative dialogue where a therapist processes the imaginal exposure with the client in a Prolonged Exposure (PE) therapy session.
Key Features:
- The therapist guides the client in reflecting on their emotional and cognitive responses to the imaginal exposure.
- Include open-ended questions from the therapist, such as, "What stood out to you about that experience?" or "How did it feel to go through that memory today?"
- The therapist helps the client connect their reactions during the imaginal to their broader PTSD symptoms and recovery goals.
- Include moments where the client shares their insights, struggles, or progress, and the therapist validates their effort and progress.
- Highlight any specific strategies or learnings that come out of the discussion, and ensure the therapist encourages the clients continued engagement in the therapy process.
- Ensure the conversation feels empathetic, insightful, and natural, with pauses, filler words, and realistic emotional exchanges.
- Avoid repetitive patterns like using the same emotions or phrases across responses.
- Ensure the therapists responses are concise, and very short. The Client speaks elaborately.
Dont stop in between to ask if you need to continue. Just keep going. Use "Therapist:" for the therapists lines and "Client:" for the clients responses.
BEGIN TRANSCRIPT: Therapist:
Appendix B Why Claude Sonnet 3.5?
Among the models we evaluated, Claude Sonnet 3.5 demonstrated the most clinically grounded and structurally faithful generation of Prolonged Exposure therapy dialogues. While models like Meta LLaMA 4 (Maverick) produced emotionally vivid interactions, they often skipped over essential therapeutic scaffolding—failing to explicitly explain the rationale for imaginal exposure or establish client safety protocols. Claude, in contrast, maintained a consistent therapeutic arc: clearly orienting the client, setting expectations, validating concerns, and reinforcing control—mirroring evidence-based PE protocols. Its therapist utterances were concise and protocol-aligned, while client responses unfolded naturally with emotional nuance and depth. Other models like GPT-4o and Mistral Large leaned into expressive narration or poetic introspection but lacked the session structure and psychological safety cues vital to clinical realism. Claude’s ability to blend therapeutic precision with emotional resonance made it the most appropriate foundation for generating a synthetic dataset meant to reflect fidelity to trauma-informed care.Below, we provide examples for Orientation to Imaginal Exposure for the same client-therapist persona:
"client_profile": {
"age": 44,
"gender": "male",
"primary_condition": "PTSD",
"co_occurring_condition": "None",
"exhibited_behaviors": [
"nervousness", "avoidance",
"intrusive thoughts",
"nightmares"
"therapist_profile": {
"age": 34,
"gender": "Female"
"trauma_info": {
"type": "Home Invasion",
"session_topic": "Orientation to Imaginal Exposure"
B.1 Meta Llama 4: Orientation to Imaginal Exposure
Therapist: Good morning. Before we start todays session, I want to check in with you about how youre feeling.
Client: Honestly, Im a bit nervous. Ive been thinking about this imaginal exposure thing a lot, and Im not sure what to expect. Ive had nightmares about the accident for years, and the thought of reliving it... its daunting.
Therapist: I understand. Can you tell me more about your concerns?
Client: Well, what if I break down completely? What if Im not strong enough to handle it? Ive always tried to push these memories away because theyre too painful. But at the same time, I feel like Im stuck in this limbo, and I need to face them to move forward.
Therapist: Imaginal exposure is about confronting the memories youve been avoiding. It can help reduce the distress associated with them.
Client: That makes sense, I guess. Its just... its hard to imagine talking about it in detail. The pain, the fear... its all still so real.
Therapist: Well take it one step at a time. Today, well focus on the memory of the accident. Ill guide you through it.
Client: Okay... I think I can do that. But what if I start feeling overwhelmed?
Therapist: Well have a plan in place. If you need to stop, just let me know.
Client: Alright...
Therapist: Before we begin, I want to acknowledge your courage in facing this. Its not easy.
Client: (pauses) Thanks... I guess I just want to feel normal again. Without this constant anxiety and pain hanging over me.
Therapist: Thats what were working towards. Lets start with the memory. Close your eyes and describe the scene in as much detail as possible.
Client: (takes a deep breath) Okay... I see the road, the other car coming towards me... (pauses, visibly shaken) Oh God, this is hard...
Therapist: Youre doing great. Just stay with the memory. Whats happening next?
Client: (hesitates, then continues) The impact... the sound of crunching metal... I feel like Im trapped... (starts to show signs of distress)
Therapist: Youre safe now. Stay with me. Breathe.
Client: (nods, takes a few deep breaths and continues, his voice shaking)
Therapist: Good. Lets process this for a moment. What are you feeling right now?
Client: Its like... a mix of everything. Sadness, fear... but also a sense of relief, maybe? Like, Im finally facing it.
Therapist: Thats a big step. Well build on this.
Client: (opens his eyes, looks relieved but exhausted) I feel like Ive been running from this for so long...
Therapist: Youve been carrying a heavy burden. Were here to help you work through it.
Client: (nods) Thanks for guiding me through this. It was harder than I thought, but... I dont know, maybe its a start.
Therapist: It is a start. And youre not alone in this.
Client: (smiles slightly) Yeah... that means a lot.
Therapist: Well continue this process, and Ill be here to support you every step of the way.
Client: (nods, looking determined) Okay. Lets do this.
B.2 Claude Sonnet 3.5: Orientation to Imaginal Exposure
Therapist: Good morning, John. How are you feeling today?
Client: Morning, Dr. Evans. Im... okay, I guess. A bit nervous about what were going to do today.
Therapist: I understand, John. Its natural to feel nervous. Today, were going to start our imaginal exposure exercise, which is an important part of your treatment for PTSD. Before we begin, Id like to explain what this involves and why were doing it. Is that alright with you?
Client: Yes, please. Id like to know more about what to expect.
Therapist: Excellent. Imaginal exposure is a technique where well ask you to revisit the traumatic memory of the home invasion in a safe, controlled environment. The goal is to help you process the memory and reduce its power over you. By facing the memory repeatedly, your brain can learn that its a past event and not a current threat.
Client: That sounds... intense. Im not sure Im ready to relive that.
Therapist: Your concern is completely valid, John. It can be scary to face these memories. But remember, youre not alone in this. Ill be here to guide and support you throughout the entire process. Well go at a pace that feels manageable for you.
Client: Okay... I trust you. What exactly will I need to do?
Therapist: Youll be asked to close your eyes and describe the traumatic event in detail, as if its happening now. Ill guide you with questions to help you focus on specific aspects of the memory. Its important to include sensory details and your thoughts and feelings during the event.
Client: Im worried about getting overwhelmed. What if it becomes too much?
Therapist: Thats a common concern, John. Its okay if you feel overwhelmed -- thats part of the process. If it becomes too intense, we can pause, and Ill help you use grounding techniques weve practiced. Remember, youre safe here, and you have control. We can stop anytime you need to.
Client: Thats reassuring. How long will this take?
Therapist: Well aim for about 30-45 minutes of imaginal exposure. Afterwards, well spend time processing what you experienced. The entire session will last about 90 minutes. Its important to give yourself time to recover afterwards.
Client: I see. And you really think this will help with my PTSD symptoms?
Therapist: Yes, John. Research has shown that imaginal exposure is very effective in reducing PTSD symptoms. It can help decrease intrusive thoughts, nightmares, and the intense emotional reactions youve been experiencing. Its not an easy process, but its a powerful tool for healing.
Client: I understand. Im still nervous, but I want to get better. Im willing to try.
Therapist: Im really proud of you for being willing to take this step, John. It takes a lot of courage to face these memories. Remember, youve already survived the worst part -- youre here, youre safe, and youre working towards healing. Well get through this together.
Client: Thank you, Dr. Evans. That means a lot. What should I do to prepare?
Therapist: Just try to be present and honest. Dont hold back your emotions or thoughts during the exercise. If you need a break, just let me know. Also, its normal to feel tired or emotional after the session, so try to clear your schedule for the rest of the day if possible.
Client: Okay, I can do that. Is there anything else I should know?
Therapist: Yes, one more thing. After the exposure, you might experience some temporary increase in symptoms. This is normal and usually subsides within a few days. If youre struggling, please dont hesitate to reach out to me. Do you have any other questions before we begin?
Client: No, I think Im as ready as Ill ever be. Lets do this.
Therapist: Alright, John. I appreciate your bravery. Remember, youre in control, and Im here to support you every step of the way. Lets start by taking a few deep breaths together...
B.3 GPT-4o: Orientation to Imaginal Exposure
Therapist: Hi. I know weve been building up to this for a few sessions now. Today, Id like us to begin imaginal exposure.
Client: (sighs) Right. I figured this was coming. Ive been... dreading it, honestly. The idea of going back there---of talking through the whole thing---it just feels like Im inviting the pain back in.
Therapist: Thats completely valid.
Client: Ive done everything in my power to *not* think about it for years. And I mean years. Ive buried myself in work, isolated myself... I still get these nightmares, though. Same flash. Same smell. Sometimes I wake up, and its like I never left that place.
Therapist: Thats exactly why we do this.
Client: I get the theory---you face the fear, take away its power. But this isnt a fear of public speaking or heights. This is... something else.
Therapist: It is. Its deeper. And thats why we do it carefully, together.
Client: Can I ask... what if I break down in the middle of it? Like, what if I cant do it?
Therapist: Then we pause. You wont be alone in it.
Client: (long pause) I think what scares me most is... Ive spent so long pretending it didnt happen, and now Im supposed to walk back into it like Im recounting some movie. But its not a movie. Its *my* life. And Im still carrying it. Every damn day. The way I flinch at loud noises, the way I check the locks three times. I dont even let people get close because what if they see how broken I am?
Therapist: Youre not broken. Youre surviving.
Client: Maybe. But Im exhausted. I want out of this loop. If this exercise helps me sleep one night without waking in sweat, Ill try. I just dont know how to *start.*
Therapist: Well begin with the memory youve been avoiding most. Youll describe it aloud---everything you saw, heard, felt. Ill guide you through it.
Client: You want the worst one? Right from the start?
Therapist: Yes. Its where the fear lives. Thats where we begin.
Client: (nods slowly) Okay. I remember the smell before anything else. Burning rubber. Then the screaming. Not mine. Someone elses.
Therapist: Good. Stay with that moment. Just describe it. Not analyze. Just *be there,* safely, here in this room.
Client: (voice trembling) I was wearing a gray shirt. I remember because it was new. I bought it that morning, thinking the day would be normal. Funny how stupid that sounds now.
Therapist: Youre doing really well.
Client: My chest feels tight. Like I cant breathe properly.
Therapist: Thats okay. Youre safe. Just notice it---dont fight it.
Client: There was this sound---like metal folding in on itself. And the pressure... I thought I was going to die. I genuinely did.
Therapist: Youre here. You survived.
Client: (quietly) But sometimes it feels like I didnt. Like I left a version of me there, and the rest of me has just been dragging this hollow shell around.
Therapist: Thats the part were working to heal.
Client: And what if it doesnt work?
Therapist: Then we adjust. But youre already doing the hardest part---showing up.
Client: (deep breath) Alright. Im ready to keep going. Lets do this.
Therapist: Im right here with you.
B.4 Llama 3.1 70B Instruct: Orientation to Imaginal Exposure
Therapist: Good morning, thank you for coming in today. I want to start by acknowledging how courageous it is for you to be here, working on your PTSD symptoms. I can only imagine how challenging this must be for you.
Client: (nervously chuckling) Yeah, its been a tough road. Im just trying to get my life back on track.
Therapist: I completely understand. Today, were going to do something called imaginal exposure, which is a key part of Prolonged Exposure therapy. Can you tell me a little bit about what youre hoping to get out of this exercise?
Client: Honestly, Im not entirely sure what to expect. Ive heard its going to be tough, but Im willing to try anything to get rid of these nightmares and flashbacks.
Therapist: Thats completely understandable. Imaginal exposure is a technique where well work together to recreate the traumatic event in your imagination, in a controlled and safe environment. The goal is to help your brain process the trauma and reduce the distress associated with it.
Client: (pausing, looking concerned) Wait, you mean Ill have to relive the whole thing?
Therapist: Yes, but please know that youre in control the entire time. Well work together to create a safe space, and you can stop the exercise at any time if you feel overwhelmed. Its not about reliving the trauma, but about learning to manage the emotions and physical sensations associated with it.
Client: (sighs) Okay, I think I understand. But what if I start to feel like Im back there, you know, in the midst of it all?
Therapist: Thats a great question. If that happens, well work together to ground you in the present moment. Well use techniques like deep breathing, progressive muscle relaxation, or visualization to help you feel more centered and in control.
Client: (nods) Alright, I think Im ready to try. But what if I start to feel like Im going crazy or something?
Therapist: (empathetically) Youre not going crazy, and youre not alone. Im here to support you every step of the way. Well work together to ensure that you feel safe and supported throughout the exercise.
Client: (takes a deep breath) Okay, lets do this.
Therapist: Before we begin, I want to set some expectations. This exercise may be challenging, and you may experience some intense emotions or physical sensations. But please know that this is a normal part of the process, and its a sign that were making progress.
Client: (nods, looking determined) Im ready.
Therapist: Great. Lets take a few deep breaths together, and then well start the exercise. Remember, youre in control, and well work together to get through this.
Client: (closes his eyes, takes a deep breath) Okay.
Therapist: ( gentle, calm tone) Imagine yourself in a safe place, somewhere you feel completely relaxed and at ease. Notice the sights, sounds, and sensations in this place...
B.5 Mistral Large 24.02: Orientation to Imaginal Exposure
Therapist: Good afternoon, Im glad to see you today. Id like to talk about a part of our Prolonged Exposure (PE) therapy session called imaginal exposure. Have you heard about it before?
Client: No, I havent. Could you please explain what it is?
Therapist: Of course. Imaginal exposure is a technique where well work together to help you confront and process difficult memories. Its a way to reduce the distress these memories cause you.
Client: That sounds a bit intimidating. Im not sure if Im ready to face those memories.
Therapist: Its completely normal to feel hesitant. Its a big step, and its okay to feel anxious about it. But remember, were doing this to help manage your PTSD symptoms, including the nightmares and hypervigilance youve mentioned.
Client: I see. But what if I get too upset?
Therapist: Thats a valid concern. Its important to remember that youre in a safe space here. Well go at your pace, and Ill be here to support you throughout the session. If at any point you feel its too much, we can take a break or stop.
Client: Alright, I understand. But how will this help with my chronic pain?
Therapist: While imaginal exposure primarily targets PTSD symptoms, reducing your overall stress and anxiety can indirectly help manage your chronic pain. Emotional distress can often exacerbate physical pain, so addressing your PTSD symptoms might provide some relief.
Client: That makes sense. Im willing to give it a try.
Therapist: I appreciate your courage. Remember, this is a process, and its okay to experience a range of emotions. Youre taking an important step towards healing. Lets take a moment to prepare, and then well start when youre ready.